Erica buys a bag of cookies that contains 6 chocolate chip cookies, 6 peanut butter cookies, 4 sugar cookies and 5 oatmeal cookies. What is the probability she selects a sugar cookie, eats it, then randomly selects an oatmeal cookie? Express the answer as a reduced fraction

rica buys a bag of cookies that contains 6 chocolate chip cookies, 6 peanut butter cookies, 4 sugar cookies and 5 oatmeal cookies. What is the probability that Erica randomly selects a sugar cookie from the bag, eats it, then randomly selects an oatmeal cookie? Express you answer as a reduced fraction. Points possible: 2 This is attempt 1 of S Submit
Answer #1

Here we are given that erica buys a bag of cookie that contains 6 choColate chips,6 peanut,4 sugar and 5 oatmeal cookie,then we are to find probability she selext a sugar cookie and after eating that oatmeal cookie

Probabilty=(4/21)*(5/20)=1/21